Drift Tube Linac (DTL)
Installation of DTL4 is ongoing in the Normal Conducting Linac. Cabling for cooling water and electrical is being connected.
DTL3 is scheduled for transport to the tunnel in October.
The Cryogenic Distribution System (CDS) for the Spoke section is subject to leak tests. The ESS vacuum group is working alongside Polish Kriosystem to repair any small leaks detected.
Proton Beam Drift Room
The 12,3-metre-long Proton Beam Drift Room pipe is in place and leak tested. Concrete plugs have now been put in place, filling up the space around the pipe.
Mock-Up and Test Stand (MUTS)
The shaft connection to the Target wheel has been vacuum tested and approved. The drive unit was assembled on top of the shaft, with electrical connections ongoing. The Target Monitoring Plug has also been reassembled.
The south grid shielding block has been placed inside the Monolith vessel. The piece weights 28 tonnes and will be followed by the equivalent block on the north side. These units have cooling pipes welded to the steel blocks.
Neutron guide optics assembly
With five Neutron Beam Port Inserts recently delivered to site, optics assembly has been completed for the FREIA insert, with leak tests following.
Next in line is the DREAM insert, to be followed by ESTIA, SKADI and VESPA.
All optics integration is performed inside a cleanroom tent.
Final assembly of the Neutron Beam Port Plugs (NBPPs) is ongoing in the E01 hall, with installation in the north sector to follow.
The installation of shielding steel walls has started for the LOKI cave. A neutron absorption material (a 5 mm thick layer of B4C, i.e. Boron carbide) has been attached to the inside of the wall panels, with the purpose to prevent backscattering of neutrons to the detectors.
Cable pulling in the under-tank raceways will follow – with parts of the walls in place.
Commissioning tests for the BIFROST bandwidth chopper continues.